Find the maximum value of resolving power of a grating 3 cm wide having 5000 lines per cm, if the wavelength of light used is 5890 A0.


1
Expert's answer
2021-06-08T09:37:14-0400

Answer:-

we know that the number of lines is 5000 lines per cm

"d=\frac{1}{5000}cm\\
d=2\times10^{-6}m\\"

now we have

"\theta=\frac{\lambda}{d}\\"

now we have resolving power given as

"RP=\frac{d}{\lambda}\\
RP=\frac{2\times10^{-6}}{5890\times 10^{-10}}"

"RP=3.4"